a banker divided $4200between two accounts, one paying 4% ANNUAL INTEREST AND THE SECOND PAYING 3.5% ANNUAL INTEREST. express THE AMOUNT INVESTED IN THE 4% ACOUNT IN TERMS OF THE AMOUNT IN THE 3.5% ACCOUNT
Answer by Theo(13342)

let x equal the amount invested in the 3.5% account.
let y equal the amount invested in the 4% account.

x + y = 4200

solve for y to get:

y = 4200 - x

that equation is the amount invested in the 4% account in terms of the amount invested in the 3.5%.

you don't know the value of x and y, and i don't believe you can find the value of x and y because you don't have enough information to determine it.

if you know the total interest, or the total overall per cent interest, then you could figure it out, but that information hasn't been provided.

i believe the overall interest rate would have to be somewhere between 3.5% and 4%.

otherwise the problem couldn't be solve.

i'll assume, for argument's sake, that it was 3.8%.

you would then have 2 equations that need to be solved simultaneosuly.

they would be x + y = 4200 and .035x +.04y = .038 * 4200

you would solve these equations simultaneously, and come up with:

x = 1680 and y = 2520.

the overall interest rate is .038 * 4200 = 159.6

.035 * 1680 + .04 * 2520 = 159.6, confirming the solution is correct.

but, you can't do that, because you don't know what the overall interest rate is, nor do you know what the overall interest is.

i'm pretty sure this is correct, unless i'm missing something.

hopefully, it's what you are looking for.

if not, give me some more information so maybe i can determine better what it is that you are looking for.